# Rehabilitation Exercise Quality Assessment through Supervised Contrastive Learning with Hard and Soft Negatives

### v1

@@ Abstract
Exercise-based rehabilitation programs have proven to be effective in enhancing the quality of life and reducing mortality and rehospitalization rates. AI-driven virtual rehabilitation, which allows patients to independently complete exercises at home, utilizes AI algorithms to analyze exercise data, providing feedback to patients and updating clinicians on their progress. These programs commonly prescribe a variety of exercise types, leading to a distinct challenge in rehabilitation exercise assessment datasets: while abundant in overall training samples, these datasets often have a limited number of samples for each individual exercise type. This disparity hampers the ability of existing approaches to train generalizable models with such a small sample size per exercise. Addressing this issue, our paper introduces a novel supervised contrastive learning framework with hard and soft negative samples that effectively utilizes the entire dataset to train a single model applicable to all exercise types. This model, with a Spatial-Temporal Graph Convolutional Network (ST-GCN) architecture, demonstrated enhanced generalizability across exercises and a decrease in overall complexity. Through extensive experiments on three publicly available rehabilitation exercise assessment datasets, the University of Idaho-Physical Rehabilitation Movement Data (UI-PRMD), IntelliRehabDS (IRDS), and KInematic assessment of MOvement and clinical scores for remote monitoring of physical REhabilitation (KIMORE), our method has shown to surpass existing methods, setting a new benchmark in rehabilitation exercise assessment accuracy.
@@
Rehabilitation Exercise, Action Quality Assessment, Supervised Contrastive Learning, Spatial-Temporal Graph Convolutional Networks
@@ 1 Introduction
Patients undergoing treatments related to cardiac, stroke, and other injuries are often referred to rehabilitation programs for swift recovery. These programs aim to enhance the quality of life of these patients, improving their ability to live independently and reducing their risk of re-hospitalization and death [1]. Typically, these programs emphasize prescribed exercises to restore mobility, muscle mass, and overall bodily strength [2, 3, 4]. Traditionally provided in a clinical setting, these programs often face long wait times, staff shortages, and logistical challenges, such as transportation and scheduling [5, 6, 7]. Virtual and home-based rehabilitation [8, 7, 9] offers a flexible alternative that overcomes these challenges and delivers benefits akin to in-clinic sessions [10, 11]. Leveraging data from virtual rehabilitation sessions, artificial intelligence (AI) can assess the quality of exercises, the patient’s recovery progression, and their risks of dropping out of the programs [12]. These approaches may utilize a variety of sensors, including wearable sensors or cameras, to monitor patients’ movements. AI algorithms analyze this data during exercises [8, 13, 14]. This analysis provides valuable feedback to patients on exercise quality and completion while enabling clinicians to track progress and personalize interventions effectively [13, 15, 14].
@@
Rehabilitation exercise quality assessment is grounded in objective measures. These include adherence to prescribed sets and repetitions of exercises [16, 17], consistency in exercise execution, ensuring proper technique and quality of movement, and maintaining correct posture of different body parts [18, 4]. Current methods in automated assessment of rehabilitation exercises mainly utilize three types of data: acceleration data from inertial wearable sensors, video data from RGB or depth cameras, and body joint data obtained either through sensors such as Kinect or extracted from RGB videos using computer vision techniques [19, 20] [4, 14]. Past studies in general human activity analysis have underscored the significance of analyzing body joints [21, 4]. In rehabilitation exercise quality assessment, this body joint analysis approach mirrors the methods clinicians use to evaluate exercise quality and technique [18]. Body joints are less affected by changes in lighting and background, making them a stable data source for analysis. This paper focuses on assessing exercises based on sequences of body joints through space and time using Spatial-Temporal Graph Convolutional Networks (ST-GCNs) [21, 22, 23, 24].
@@
Patients in rehabilitation programs receive group and individual level exercises that are tailored to their specific needs, stage of their rehabilitation program, age, sex, and comorbidity [3]. Examples of rehabilitation exercise types include standing shoulder abduction, right elbow flexion, and deep squats [25, 26, 18]. A primary challenge in current methods of rehabilitation exercise quality assessment [21, 22, 23, 24] lies in their dependency on distinct models for each type of rehabilitation exercise. Training individual models for each exercise type is problematic. Existing datasets, such as the University of Idaho-Physical Rehabilitation Movement Data (UI-PRMD) [25], often have a high total number of samples that are spread sparsely across various exercise types, resulting in insufficient samples for each specific exercise type. This poses a challenge for training exercise-type-specific deep neural networks that require large amounts of data [4, 14, 27].
@@
This paper aims to address the above-mentioned issue by proposing a unified model that leverages all samples across different exercise types in a dataset, rather than separate models limited to their respective exercise samples. This approach considers that each exercise type falls within a specific range of acceptable latent spaces for correct assessment, with deviations indicating incorrectness. This principle is uniformly applicable across all exercise types using a single exercise assessment model. To this end, this paper makes two key contributions: (i) introducing a novel supervised contrastive learning method [28] equipped with hard [29] and soft negatives specifically designed for rehabilitation exercise quality assessment, where a single ST-GCN model abstracts the general assessment process in a dataset, and (ii) demonstrating quantitative superiority over previous methods, enhancing the state-of-the-art in rehabilitation exercise quality assessment on three public datasets, UI-PRMD [25], IntelliRehabDS (IRDS) [26], and KInematic assessment of MOvement and clinical scores for remote monitoring of physical REhabilitation (KIMORE) [18].
@@ 2 Related Work
This section reviews existing approaches, including both general deep learning techniques (non-ST-GCN) and ST-GCN-based methods, for rehabilitation exercise quality assessment based on body joints [4, 14]. With regard to the focus and contributions of the method introduced in this paper, this section explores strategies previously employed to address the specific settings of existing datasets for rehabilitation exercise quality assessment. These settings typically feature a variety of exercise types with a limited number of samples for each type [4, 14, 25, 26, 18]. The majority of existing methodologies, as outlined below, have involved the development of exercise-type-specific models trained on samples from specific exercise types in datasets [4, 30, 24, 23].
@@ 2.2 ST-GCN-based Methods
This paper introduces a novel rehabilitation exercise quality assessment approach using supervised contrastive learning [28] with hard [29] and soft negatives for training a general, non-exercise-type-specific ST-GCN [21, 42] model. This method benefits from leveraging training samples encompassing all exercise types in the complete dataset.
@@ 3.2 Supervised Contrastive Learning with Hard and Soft Negatives
$\begin{split}\mathcal{L}&=\sum_{i\in\beta^{+}}\frac{-1}{\|\beta^{+}_{y_{i}}\|}\phantom{=}\sum_{j\in\beta^{+}_{y_{i}},\,j\neq i}\log\frac{\exp(\mathrm{sim}(\tilde{\mathbf{v}}_{i},\tilde{\mathbf{v}}_{j})/\tau)}{\sum_{k\in\beta^{+},\,k\neq i}\exp(\mathrm{sim}(\tilde{\mathbf{v}}_{i},\tilde{\mathbf{v}}_{k})/\tau)+K}\end{split}$ (2)
$K=\sum_{\ell\in\beta^{-}}\exp(\mathrm{sim}(\tilde{\mathbf{v}}_{i},\tilde{\mathbf{v}}_{\ell})/\tau),$ (3)
@@ 3.3 Inferencing
$\mathcal{D}$ is defined as the index partition for dataset $D$, identically to the mini-batch-wise index partitions in subsection 3.2. For a given arbitrary exercise type $c$, a reference representation is generated by averaging the representations of the correct samples for that exercise type, computed as follows:
$\mathbf{r}_{c}=\sum_{\ell\in\mathcal{D}^{+}_{c}}\frac{g(f(x_{\ell}))}{\|\,g(f(x_{\ell}))\|_{2}}.$ (5)
Making inference for a sample with index $i$ is performed as follows:
@@ 4.2 Datasets
    KIMORE [18] is a rehabilitation exercise dataset consisting of 5 exercises. It constitutes healthy subjects and patients with motor dysfunctions. Medical professionals assessed all performances with a clinical score for correctness, ranging from 0 to 50. Body joints data was collected with the Kinect One sensor at 30 fps. This dataset presents a significant challenge due to its regressive nature, compounded by the limited amount of data available, with only  70 samples per exercise type across all participants. The 5 types of exercises are lifting of arms (k01), trunk lateral tilt (k02), trunk rotation (k03), pelvis rotation (k04), and squatting (k05).
@@ 4.3 Experimental Setting
In the datasets for rehabilitation exercise quality assessment [25, 26, 18], each type of exercise specifies a series of essential body-joint movements to assess the exercise’s correctness. These movements are characterized by the joints’ trajectory through space and time, enabling the conceptualization of each movement through a sequence of high-order spatial-temporal embeddings. Considering ST-GCN as the optimal model for learning such embeddings [24, 23, 37, 38, 22], the encoder’s architecture in the proposed method employs an 8-layer ST-GCN as elaborated by Yan et al. [21]. All temporal layers of ST-GCN blocks are downsized using the ResNet bottleneck architecture [42]. A feed-forward layer serves as the projection head, transforming embeddings from a dimensionality of 256 to 128.
@@ 4.4 Experimental Results
As explained in section 2, one of the recent pioneering works on rehabilitation exercise quality assessment is by Zheng et al. [23], wherein ST-GCNs were equipped with a RI descriptor to stabilize models against rotational variations in body-joint data. To ensure a fair comparison with relevant works, identical 3:1 training-validation set split as in Zheng et al. [23] were adopted. Table 1 depicts the accuracy of the proposed method compared to previous relevant methods on individual exercises and on average on the UI-PRMD [25] and IRDS [26] datasets. In addition to the work by Zheng et al. [23], denoted as ST-GCN with RI in Table 1, the results of two previous non-ST-GCN methods on body-joint-based action analysis using LSTMs [51] and CNNs [52] are also reported [23]. The results of the proposed method equipped with the RI descriptor introduced by Zheng et al. [23] are also presented in Table 1. According to Table 1, on UI-PRMD [25], the proposed method surpassed previous methods in exercises u02, u03, and on average. In nine out of ten exercises, the proposed method achieved an accuracy of 1. On IRDS [26], a dataset with imbalanced distributions of samples across correct and incorrect exercises, the proposed method’s superiority was more evident, specifically in exercises i01-i04, i06, and i08. The integration of RI [23] with the proposed method either mirrored the results of the proposed method or provided a slight boost in accuracy.
@@
To explore the effectiveness of retaining the projection head during inference, as discussed in subsection 3.3, the accuracy of the proposed method equipped with the RI descriptor with and without the projection head on the UI-PRMD dataset [25], was evaluated using five-fold cross-validation and is reported in Table 4. In all ten exercises of UI-PRMD [25], keeping the projection head for inference led to an accuracy improvement. A previous study on driver anomaly detection which kept the projection head within a contrastive learning setting has found similar results [45]. This improvement was more pronounced in exercises that are uni-lateral or vertically asymmetrical, such as inline lunge (u03) or straight leg raise (u06). In UI-PRMD, the performance of subjects in exercises depends on their dominant side. The single ST-GCN struggles in this scenario because representations for left and right leg raises will be spatially different. This spatial divergence leads to instability in the reference representation, resulting in sub-optimal evaluations. Therefore, projecting side-variant embeddings into an equivalent space is crucial. To reemphasize from the reverse perspective, as shown in Table 4, ST-GCN-only representations relatively suffice for symmetrical exercises such as deep squats (u01) and sit to stand (u05).
Table 4: Exploring the importance of keeping the projection head during inference making. Accuracy of the proposed method with and without projection head on u01-u10 in the UI-PRMD [25] dataset through five-fold cross-validation is reported. Bolded values denote the best results.

### v2

@@ Abstract
Exercise-based rehabilitation programs have proven to be effective in enhancing the quality of life and reducing mortality and rehospitalization rates. AI-driven virtual rehabilitation, which allows patients to independently complete exercises at home, utilizes AI algorithms to analyze exercise data, providing feedback to patients and updating clinicians on their progress. These programs commonly prescribe a variety of exercise types, leading to a distinct challenge in rehabilitation exercise assessment datasets: while abundant in overall training samples, these datasets often have a limited number of samples for each individual exercise type. This disparity hampers the ability of existing approaches to train generalizable models with such a small sample size per exercise type. Addressing this issue, this paper introduces a novel supervised contrastive learning framework with hard and soft negative samples that effectively utilizes the entire dataset to train a single model applicable to all exercise types. This model, with a Spatial-Temporal Graph Convolutional Network (ST-GCN) architecture, demonstrated enhanced generalizability across exercises and a decrease in overall complexity. Through extensive experiments on three publicly available rehabilitation exercise assessment datasets, UI-PRMD, IRDS, and KIMORE, our method has proven to surpass existing methods, setting a new benchmark in rehabilitation exercise quality assessment.
@@
Rehabilitation Exercise, Action Quality Assessment, Supervised Contrastive Learning, Graph Convolutional Networks, Hard and Soft Negatives
@@ 1 Introduction
Patients undergoing treatments related to cardiac, stroke, and other injuries are often referred to rehabilitation programs for swift recovery. These programs aim to enhance the quality of life of these patients, improving their ability to live independently and reducing their risk of re-hospitalization and death [1]. Typically, these programs emphasize prescribed exercises to restore mobility, muscle mass, and overall bodily strength [2, 3, 4]. Traditionally provided in a clinical setting, these programs often face long wait times, staff shortages, and logistical challenges, such as transportation and scheduling [5, 6, 7]. Virtual and home-based rehabilitation [8, 7, 9] offers a flexible alternative that overcomes these challenges and delivers benefits akin to in-clinic sessions [10, 11]. Leveraging data from virtual rehabilitation sessions, Artificial Intelligence (AI) can assess the quality of exercises, the patient’s recovery progression, and their risks of dropping out of the programs [12]. These approaches may utilize a variety of sensors, including wearable sensors or cameras, to monitor patients’ movements. AI algorithms analyze this data during exercises [8, 13, 14]. This analysis provides valuable feedback to patients on exercise quality and completion while enabling clinicians to track progress and personalize interventions effectively [13, 15, 14, 12].
Rehabilitation exercise quality assessment is grounded in objective measures. These include adherence to prescribed sets and repetitions of exercises [16, 17], consistency in exercise execution, ensuring proper technique and quality of movement, and maintaining correct posture of different body parts [18, 4, 19, 20]. The development of AI models for rehabilitation exercise quality assessment relies on annotated datasets [21, 22, 23, 24, 4]. Experts in the field, such as rehabilitation clinicians and physiotherapists, observe patients as they complete exercises and annotate them [19, 18]. The exercise data and their corresponding annotations are then used for the development of AI models. In some datasets, such as the KInematic assessment of MOvement and clinical scores for remote monitoring of physical REhabilitation (KIMORE) [18], clinically validated tools, such as the Exercise Accuracy Assessment Questionnaire (EAAQ) [20], were used for annotation resulting in real-valued numbers representing exercise quality scores. However, other datasets , such as the University of Idaho-Physical Rehabilitation Movement Data (UI-PRMD) [25] and IntelliRehabDS (IRDS) [26], sufficed with annotating rehabilitation exercises as correct or incorrect binary values. This inconsistency [27] hinders the development of AI models applicable across datasets. A model trained on IRDS for a binary classification problem is not directly applicable to KIMORE, which requires solving a regression problem.
@@
Current methods in automated rehabilitation exercise quality assessment mainly utilize three types of data: acceleration data from inertial wearable sensors, video data from RGB or depth cameras, and body joint data obtained either through sensors such as Kinect or extracted from RGB videos using computer vision techniques [28, 29, 17] [4, 14]. Past studies in general human activity analysis have underscored the significance of analyzing body joints [21, 4]. In rehabilitation exercise quality assessment, this body joint analysis approach mirrors the methods clinicians use to evaluate exercise quality and technique [18]. Body joints are less affected by changes in lighting and background, making them a stable data source for analysis. This paper focuses on assessing exercises based on sequences of body joints through space and time using Spatial-Temporal Graph Convolutional Networks (ST-GCNs) [21, 22, 23, 24].
@@
Patients in rehabilitation programs receive group and individual level exercises that are tailored to their specific needs, stage of their rehabilitation program, age, sex, and comorbidity [3]. Examples of rehabilitation exercise types include standing shoulder abduction, right elbow flexion, and deep squats [25, 26, 18]. A primary challenge in current methods of rehabilitation exercise quality assessment [21, 22, 23, 24] lies in their dependency on distinct models for each type of rehabilitation exercise. Training individual models for each exercise type is problematic. Existing datasets, such as UI-PRMD [25], often have a high total number of samples that are spread sparsely across various exercise types, resulting in insufficient samples for each specific exercise type. This poses a challenge for training exercise-type-specific deep neural networks that require large amounts of data [4, 14, 30].
@@
This paper aims to address the above-mentioned issue by proposing a unified model that leverages all samples across different exercise types in a dataset, rather than separate models limited to their respective exercise samples. This approach considers that each exercise type falls within a specific range of acceptable latent spaces for correct assessment, with deviations indicating incorrectness. This principle is uniformly applicable across all exercise types using a single exercise assessment model. To this end, this paper makes two key contributions: (i) introducing a novel supervised contrastive learning method [31] equipped with hard [32] and soft negatives specifically designed for rehabilitation exercise quality assessment, where a single ST-GCN model abstracts the general assessment process in a dataset, and (ii) demonstrating quantitative superiority over previous methods, enhancing the state-of-the-art in rehabilitation exercise quality assessment on three public datasets, UI-PRMD [25], IRDS [26], and KIMORE [18]. In an attempt to develop an exercise quality assessment model applicable across datasets, a model was trained on IRDS [26] through the proposed contrastive learning approach. This model then underwent transfer learning to make inferences on KIMORE [18] as a regression problem, showing advancements compared to previous works.
@@ 2 Related Work
This section reviews existing approaches for rehabilitation exercise quality assessment, including both general deep learning techniques (non-ST-GCN) and ST-GCN-based methods specifically designed for assessing rehabilitation exercise quality based on body joints [4, 14]. With regard to the focus and contributions of the method introduced in this paper, this section explores strategies previously employed to address the specific settings of existing datasets for rehabilitation exercise quality assessment. These settings typically feature a variety of exercise types with a limited number of samples for each type [4, 14, 25, 26, 18].
@@ 2.2 ST-GCN-based Methods
The majority of existing methodologies, as outlined above, have involved the development of exercise-type-specific models trained on samples from specific exercise types in a dataset [4, 33, 24, 23]. In contrast to the current literature, this paper introduces a novel method that leverages training samples encompassing all exercise types in a dataset, resulting in improved rehabilitation exercise quality assessment.
@@ 3.1 Background
Figure 1: (a) Variety of samples in a rehabilitation exercise training mini-batch, featuring different exercise types where each sample may be correct or incorrect, with the leftmost sample designated as the anchor. (b) From left to right, a positive sample pair and its corresponding hard negative sample pair and two soft negative sample pairs.
@@ 3.2 Supervised Contrastive Learning with Hard and Soft Negatives
$\begin{split}\mathcal{L}&=\sum_{i\in\beta^{+}}\frac{-1}{\|\beta^{+}_{y_{i}}\|}\\ &\phantom{=}\sum_{j\in\beta^{+}_{y_{i}},\,j\neq i}\log\frac{\exp(\mathrm{sim}(\tilde{\mathbf{v}}_{i},\tilde{\mathbf{v}}_{j})/\tau)}{\sum_{k\in\beta^{-}_{y_{i}}}\exp(\mathrm{sim}(\tilde{\mathbf{v}}_{i},\tilde{\mathbf{v}}_{k})/\tau)+\sum_{\ell\neq i}\exp(\mathrm{sim}(\tilde{\mathbf{v}}_{i},\tilde{\mathbf{v}}_{\ell})/\tau)}\end{split}$ (2)
@@
The numerator in the contrastive loss function in Equation 2 pertains to positive pairs, which are two correct samples of the same exercise type. The first summation in the denominator pertains to hard negative pairs, consisting of a correct and an incorrect sample of the same exercise type. The second summation in the denominator pertains to soft negative pairs, which include two correct samples of different exercise types and a correct and an incorrect sample of different exercise types. Please refer to Fig. 1.
@@
Figure 2: (a) Using all training exercise samples with the mini-batches as described in Fig. 1 and the supervised contrastive loss function in Equation 2, the spatial-temporal graph convolutional network encoder $f(\cdot)$ and fully-connected projection head $g(\cdot)$ are trained. (b) Trained $f(\cdot)$ and $g(\cdot)$ are used to generate the learned representations for all correct type $c$ training exercise samples. Weighted averaging of the learned representations results in an exercise-type-specific reference representation for type $c$ exercise. (c) Inference making by calculating the similarity between the learned representation of a test exercise sample of type $c$ with the reference representation for type $c$ exercise.
@@ 3.3 Inferencing
$\mathcal{D}$ is defined as the index partition for dataset $D$, identically to the mini-batch-wise index partitions in subsection 3.2. For a given arbitrary exercise type $c$, a reference representation is generated through a weighted average of the representations of the correct samples for that exercise type. Weight vector $\mathbf{w}$, is computed according to the inverse variance of each feature dimension from the embedding matrix $\mathbf{v_{\ell}}=g(f(x_{\ell})),\forall\ell\in\mathcal{D}^{+}_{c}$,
$\mathbf{w}=\frac{1}{\mathrm{Var}(\textbf{V})}\phantom{=}\phantom{=}\phantom{=}\phantom{=}\mathbf{r}_{c}=\frac{\mathbf{w}}{\ \|\mathbf{w}\|_{1}}\otimes\sum_{\ell\in\mathcal{D}^{+}_{c}}\mathbf{v}_{\ell}$ (4)
where $\otimes$ denotes Hadamard product. Making inference for a sample with index $i$ is performed as follows:
@@ 4.2 Datasets
    KIMORE [18] is a rehabilitation exercise dataset consisting of 5 exercises. It constitutes healthy subjects and patients with motor dysfunctions. Medical professionals assessed all performances with a clinical score for correctness [20], ranging from 0 to 50. Body joints data was collected with the Kinect One sensor at 30 fps. This dataset presents a significant challenge due to its regressive nature, compounded by the limited amount of data available, with only  70 samples per exercise type across all participants. The 5 types of exercises are lifting of arms (k01), trunk lateral tilt (k02), trunk rotation (k03), pelvis rotation (k04), and squatting (k05).
@@ 4.3 Experimental Setting
In the datasets for rehabilitation exercise quality assessment [25, 26, 18], each type of exercise specifies a series of essential body-joint movements to assess the exercise’s correctness. These movements are characterized by the joints’ trajectory through space and time, enabling the conceptualization of each movement through a sequence of high-order spatial-temporal embeddings. Considering ST-GCN as the optimal model for learning such embeddings [24, 23, 40, 41, 22], the encoder’s architecture in the proposed method employs an 8-layer ST-GCN as elaborated by Yan et al. [21]. All temporal layers of ST-GCN blocks are downsized using the ResNet bottleneck architecture [48]. A fully-connected layer serves as the projection head, transforming embeddings from a dimensionality of 256 to 128.
@@ 4.4 Experimental Results
4.4.1 Comparison with Previous Works
As explained in section 2.2, one of the recent pioneering works on rehabilitation exercise quality assessment is by Zheng et al. [23], wherein ST-GCNs were equipped with a RI descriptor to stabilize models against rotational variations in body-joint data. To ensure a fair comparison with relevant works, identical 3:1 training-validation set split as in Zheng et al. [23] was adopted. Table 1 depicts the accuracy of the proposed method compared to previous relevant methods on individual exercises and on average on the UI-PRMD [25] and IRDS [26] datasets. In addition to the work by Zheng et al. [23], denoted as ST-GCN with RI in Table 1, the results of two previous non-ST-GCN methods on body-joint-based action analysis using LSTMs [53] and CNNs [54] are also reported [23]. The results of the proposed method equipped with the RI descriptor introduced by Zheng et al. [23] are also presented in Table 1. According to Table 1, on UI-PRMD [25], the proposed method surpassed previous methods in exercises u02, u03, and on average. In nine out of ten exercises, the proposed method achieved an accuracy of 1. On IRDS [26], a dataset with imbalanced distributions of samples across correct and incorrect exercises, the proposed method’s superiority was more evident, specifically in exercises i01-i04, i06, and i08. The integration of RI [23] with the proposed method either mirrored the results of the proposed method or provided a slight boost in accuracy.
@@ 4.4.1 Comparison with Previous Works
4.4.2 Impact of Retaining the Projection Head
To explore the effectiveness of retaining the projection head during inference, as discussed in subsection 3.3, the performance of the proposed method equipped with the RI descriptor with and without the projection head on the UI-PRMD dataset [25], was evaluated using five-fold cross-validation and is reported in Table 4.
Considering the first two rows in Table 4, in all ten exercises of UI-PRMD [25], retaining the projection head for inference led to an accuracy improvement. This improvement was more pronounced in exercises that are uni-lateral or vertically asymmetrical, such as inline lunge (u03) or straight leg raise (u06). In UI-PRMD, the performance of subjects in exercises depends on their dominant side. The single ST-GCN struggles in this scenario because representations for left and right leg raises will be spatially different. This spatial divergence leads to instability in the reference representation, resulting in sub-optimal evaluations. Therefore, projecting side-variant embeddings into an equivalent space is crucial. To reemphasize from the reverse perspective, as shown in Table 4, encoder-only representations relatively suffice for symmetrical exercises such as deep squats (u01) and sit-to-stand (u05).
A previous study on driver anomaly detection which kept the projection head within a supervised contrastive learning setting found similar results [46, 47]. The superior inference results achieved while retaining the projection head can be attributed to the nature of the problem, which involves supervised contrastive learning. The idea of discarding the projection head during inference originated in self-supervised contrastive learning [45], where labels are not available. However, in supervised contrastive learning, where additional information, i.e., labels, is available, the projection head enhances the model’s ability to learn more effective representations by adding complexity.
The second and third rows in Table 4 compare the proposed method with that of Zheng et al. [23] using five-fold cross-validation setting. This is different from the comparison in Table 1, which followed a 3:1 training-validation set split, as in Zheng et al. [23]. The key differences between Zheng et al. [23] and the proposed method are twofold: (1) Zheng et al. [23] uses 10 distinct ST-GCN-based models, each trained for a specific exercise type. In contrast, the proposed method employs a single ST-GCN-based model that handles all exercise types. (2) The architecture used by Zheng et al. [23] consists of an ST-GCN model followed by fully connected layers for classification, without utilizing contrastive learning for model training. On the other hand, the proposed method incorporates contrastive learning with hard and soft negatives. It uses an ST-GCN model as the encoder within the contrastive learning framework, followed by a fully connected network as the projection head. The results show that the proposed contrastive learning-based method, specifically when the projection head is retained, is superior, achieving equal or better accuracy in seven out of ten exercise types.
To further investigate the efficacy of the representations of the encoder and projection head learned through the proposed supervised contrastive learning approach, Support Vector Machines (SVMs) with a Radial Basis Function (RBF) kernel were trained on these representations for exercise quality assessment. For the RBF kernel of the SVMs, the parameters C, and gamma were set to $1$, and $1/128=0.0078$, respectively. As shown in the last row of Table 4, promising results were obtained for U01, U02, and U04, demonstrating the effectiveness of the learned representations through the proposed method; even an SVM applied to these representations can successfully perform exercise quality assessment.
Table 4: Exploring the efficacy of learning representations through the proposed supervised contrastive learning approach and the importance of retaining the projection head during inference making. Accuracy of different approaches on u01-u10 in the UI-PRMD [25] dataset through five-fold cross-validation is reported. Bolded values denote the best results.